OpenStudy (anonymous):

Solve for the variable in this equation. 5x/3+2/7=x/9-4/7

OpenStudy (zepp):

Like this? http://puu.sh/qrBw

OpenStudy (anonymous):

Yes @zepp thanks:) Can you help me figure it out?

OpenStudy (zepp):

Yes, first, try to pass the 2/y to the right

OpenStudy (anonymous):

Im not sure i understand

OpenStudy (zepp):

Try to take out the 2/y

OpenStudy (anonymous):

There is no Y

OpenStudy (zepp):

2/7.. sorry :x

OpenStudy (anonymous):

Oh so you want me to take 2/7 out, but how do i do that?

OpenStudy (zepp):

Like this: http://puu.sh/qrI5 Sorry for the waiting time, my internet's lagging :(

OpenStudy (anonymous):

Ok now what, and its ok my internet does that too

OpenStudy (radar):

Now your equation looks like this: 5x/3 =x/9 - 6/7 after you combine the -2/7 and -4/7. Do you agree?

OpenStudy (anonymous):

Yes:)

OpenStudy (radar):

Now subtract x/9 from both sides of the equal sign getting: 5x/3 - x/9 =-6/7 did you follow that little maneuver?

OpenStudy (anonymous):

Yes i think so

OpenStudy (radar):

O.K Now multiply all terms by 9 getting 15x-x=-54/7 combine the similar terms on the left. 14x=-54/7

OpenStudy (anonymous):

Yeah ok so then i

OpenStudy (zepp):

http://puu.sh/qrNj To resume everything. :)

OpenStudy (radar):

Now you can cross multiply.